# Science Park 904, 1098 XH, Amsterdam, The Netherlands

### v1

@@ 4.3 Relation Between Outer Entropy in Two Frames
$S^{\text{(outer)}}_{E}[\mu]=\frac{\text{Area}[\mu]}{4G}.$ (95)
@@
$S^{\text{(outer)}}_{E}[\mu]=S_{f}^{\text{(outer)}}[\mu]=\frac{\text{Area}[\mu]}{4G}=\frac{\int_{\mu}f^{\prime}(R)}{4G}.$ (96)
@@
here $c$ is a positive constant, and this shows that $\Theta_{v}=0$. And the minimar condition of $\mu$ in the Einstein frame implies that $\mu$ satisfies the generalized minimar condition  [^7] It can be shown that $\nabla_{l_{E}}\theta^{E}_{k_{E}}|_{\mu}\propto\frac{f^{\prime}(R)\partial_{u}\Theta_{v}-\Theta_{v}\partial_{u}f^{\prime}(R)}{f^{\prime}(R)^{2}}|_{\mu}=\frac{\partial_{u}\Theta_{v}}{f^{\prime}(R)}|_{\mu}\leq 0$. Since $f^{\prime}(R)>0$, this implies $\nabla_{l}\Theta_{k}=\partial_{u}\Theta_{v}\leq 0$. (see section 1.1) in the $f(R)$ frame. Therefore, $\mu$ is just the generalized minimar surface in the $f(R)$ frame, and we finally show that the outer entropy of the generalized marginally trapped surface in $f(R)$ gravity is just the Wald entropy of it
@@ 5 Deriving the Results in the $f(R)$ Frame
5.1 Raychaudhuri Equation For the Generalized Expansion
@@
We define (107) as the generalized Raychaudhuri equation.
@@
Now consider the generalized marginally trapped surface $\mu$, let $N_{\pm k}(\mu)$ be the null congruence generated from $\mu$ by null geodesics $k^{a}$ and $k^{a}$ in the $f(R)$ frame. Defining $\bar{\mu}$ as the representative of $\mu$ on Cauchy slice $\Sigma$, following the definition in section 3.1. Assuming $\bar{\mu}$ lies on $N_{k}$, since $\Theta_{v}|_{\mu}=0$ and the generalized Raychaudhuri equation dictates $\partial_{v}\Theta_{v}\leq 0$, it follows that
@@ 5.2 Construction of Stationary Null Hypersurface For the Generalized Expansion
Similar to section 3, constructing a spacetime with a fixed $O_{W}^{f}[\mu]$ that maximizes $S_{\rm vN}(\rho_{B}^{f})$ requires gluing a stationary null hypersurface $N_{-k}$ for the generalized expansion, subject to specific initial data on $N_{-k}$. This will show that there exists a surface $X$ with zero generalized expansions of two null directions (the surface that we calculate the von Neumann entropy in the $f(R)$ frame), the Wald entropy of this surface is equal to the Wald entropy of $\mu$. The spacetime is then completed via a CPT reflection across $X$, analogous to the procedure in section 3 (see Figure 3).
@@ 5.3 Construction of the Generalized Extremal Surface and the Outer Entropy in the $f(R)$ Frame
Similarly to section 3.2, we complete our construction by a CPT-reflection across $X$ that takes $v\to-v$, $u\to-u$, $x^{i}\to x^{i}$. Under the CPT transformation, $\Phi$, $f^{\prime}(R)$ and $\chi_{i}$ are even, whereas $\Theta_{u}$, $\Theta_{v}$ and $\partial_{v}f^{\prime}(R)$ are odd. All quantities that are odd under CPT vanish on $X$ by construction. Therefore, the CPT-conjugate data satisfies the requisite junction conditions. The results is a second boundary $\tilde{B}$, $B$ and $\tilde{B}$ are connected by a Cauchy slice with three null segments, $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\cup N^{\prime}_{-l}[\tilde{\mu}$], we are using prime to represent CPT-conjugated submanifolds. We have now specified all data necessary to uniquely evolve characteristic initial data via the EOM of $f(R)$ gravity. The resulting spacetime $(\mathcal{M}^{\prime},g)$ has a generalized minimal surface $\mu$ with a fixed outer wedge $O_{W}^{f}[\mu]$. This spacetime contains a generalized extremal surface $X$ on the boundary of the inner wedge $I_{W}^{f}[\mu]$, which is homologous to $\mu$ and therefore to the boundary $B$.
Finally, we need to show that this generalized extremal surface $X$ has the least Wald entropy. Given the duality established in (76), $X$ must have minimal area in the Einstein frame. Therefore, in the $f(R)$ frame, $X$ should have the least Wald entropy if there exists other generalized extremal surfaces. If the constructed spacetime $(\mathcal{M}^{\prime},g^{\prime})$ contains no other generalized extremal surfaces, the proof is complete. If there exists other generalized extremal surfaces we denote it by $X^{\prime}$, define $\overline{X^{\prime}}[\Sigma]$ as the representative of $X^{\prime}$ on Cauchy slice $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\cup N^{\prime}_{-l}[\tilde{\mu}$]. There are in total three cases; If $\overline{X^{\prime}}[\Sigma]$ lies entirely on $N_{-k}$, then, because $N_{-k}$ is stationary with respect to the generalized expansion, we have
@@ 7 Conclusion
A key ingredient of our analysis is a careful treatment of conformal frames and their holographic interpretation. We first derive holographic description in the Einstein frame, where the theory can be written as Einstein gravity coupled to an auxiliary scalar degree of freedom, and where standard holographic techniques apply directly. Within this setup, we then establish an explicit correspondence between the boundary von Neumann entropy computed in the Einstein frame description and the corresponding entropy in the $f(R)$ frame. Concretely, we track how the reduced density matrix and its entropy transform under the field redefinitions and Weyl rescalings relating the two frames, and we identify the precise map that preserves the physical content of entanglement. Using this entropy dictionary, we translate the outer-entropy construction between frames, obtaining a robust outer-entropy correspondence that allows us to import the Einstein-frame maximization logic and re-express it in purely geometric language in the $f(R)$ frame. This provides the conceptual bridge needed to equate the coarse-grained outer entropy with the Wald functional of the generalized marginally trapped surface.
@@
Building on this foundation, we proceed to derive the outer entropy directly in $f(R)$ gravity. We prove a focusing theorem for an appropriately defined generalized expansion, which serves as a higher-curvature generalization of the Raychaudhuri equation and the usual null convergence/energy conditions. With this theorem in hand, we construct a stationary null hypersurface characterized by vanishing generalized expansion, and we formulate the necessary junction conditions for consistently stitching geometries across a hypersurface within $f(R)$ gravity. These tools allow us to explicitly engineer a bulk spacetime whose boundary von Neumann entropy saturates the upper bound implied by the outer-entropy maximization—i.e., we realize the extremal configuration achieving the coarse-grained maximum compatible with the fixed exterior. Finally, using the $f(R)$ focusing theorem, we show that the simple entropy defined on the boundary is indeed the holographic dual of the outer entropy in $f(R)$ gravity, thereby extending the outer entropy and simple entropy correspondence to this higher-curvature setting and show the second law of the two entropy.
@@
Looking forward, our results suggest several promising future directions. Since our construction of the coarse graining entropy is non-perturbative in classical gravity (at least to all orders in perturbation theory around equilibrium), it suggests that the non-perturbative construction of the dynamical black hole entropy is highly related to the generalized marginally trapped surface (generalized minimar surface). In a recent paper Ashtekar:2025qqa it was suggested that quasi-local horizon (QLH) is the right geometric concept for far from equilibrium black hole thermodynamics in Einstein gravity. We hope that similar concepts will appear in general higher curvature gravity. Recently there are some discussions about higher curvature focusing theorem near dynamic perturbative killing horizon Yan:2024gbz ; Yan:2025vdp , our construction of a non-perturbative focusing theorem in $f(R)$ gravity suggests similar theorem may exists in more general higher curvature gravity, for example, in $f(\text{Riemann})$ gravity. This will help us define HRT surface in general higher curvature gravity, and explain why the Wall entropy and Dong entropy Dong:2013qoa ; Wall:2015raa are the same but with different physical original.
@@ Acknowledgements.
We would like to thank Antony J. Speranza, Zihan Yan and Jiyu Cheng for discussions. In particular, we are greatly grateful to Antony J. Speranza for very patient discussions and useful comments on the draft. We are also grateful for the beautiful environment and academic atmosphere of University of Amsterdam, where this work was completed.
@@ Appendix A Cases of $f^{\prime\prime}(R)=0$
Since $g^{uu}=g^{ui}=g^{vi}=0$ on $N_{-k}$, we only need to consider $T_{uv}$ and $T_{ij}$. Imposing the initial free data conditions (129) and (128) yields $\partial_{v}T=0$. Therefore, this will imply $\partial_{v}R|_{V_{I^{j}_{R}}}=0$. In conclusion, even if there exist subregions of $N_{-k}$ where $f^{\prime\prime}(R)=0$, the result $\partial_{v}R=0$ remains valid.

### v2

@@ 4.3 Relation Between Outer Entropy in Two Frames
$S^{\text{(outer)}}_{E}[\mu]=\frac{\text{Area}_{E}[\mu]}{4G}.$ (95)
@@
$S^{\text{(outer)}}_{E}[\mu]=S_{f}^{\text{(outer)}}[\mu]=\frac{\text{Area}_{E}[\mu]}{4G}=\frac{\int_{\mu}f^{\prime}(R)}{4G}.$ (96)
@@
here $c$ is a positive constant, and this shows that $\Theta_{v}=0$. And the minimar condition of $\mu$ in the Einstein frame implies that $\mu$ satisfies the generalized minimar condition  [^7] It can be shown that $\nabla_{l_{E}}\theta^{E}_{k_{E}}|_{\mu}\propto\frac{f^{\prime}(R)\partial_{u}\Theta_{v}-\Theta_{v}\partial_{u}f^{\prime}(R)}{f^{\prime}(R)^{2}}|_{\mu}=\frac{\partial_{u}\Theta_{v}}{f^{\prime}(R)}|_{\mu}\leq 0$. Since $f^{\prime}(R)>0$, this implies $\nabla_{l}\Theta_{k}=\partial_{u}\Theta_{v}\leq 0$. (see section 1.1) in the $f(R)$ frame. Therefore, $\mu$ is just the generalized minimar surface in the $f(R)$ frame, and we finally show that the outer entropy of the generalized marginally trapped surface with generalized minimar condition in $f(R)$ gravity is just the Wald entropy of it
@@ 5 Deriving the Results in the $f(R)$ Frame
5.1 Nonlinear Raychaudhuri Equation For the Generalized Expansion
@@
We define (107) as the nonlinear Raychaudhuri equation.
@@
Now consider the generalized marginally trapped surface $\mu$, let $N_{\pm k}(\mu)$ be the null congruence generated from $\mu$ by null geodesics $k^{a}$ and $k^{a}$ in the $f(R)$ frame. Defining $\bar{\mu}$ as the representative of $\mu$ on Cauchy slice $\Sigma$, following the definition in section 3.1. Assuming $\bar{\mu}$ lies on $N_{k}$, since $\Theta_{v}|_{\mu}=0$ and the nonlinear Raychaudhuri equation dictates $\partial_{v}\Theta_{v}\leq 0$, it follows that
@@ 5.2 Construction of Stationary Null Hypersurface For the Generalized Expansion
Similar to section 3, constructing a spacetime with a fixed $O_{W}^{f}[\mu]$ that maximizes $S_{\rm vN}(\rho_{B}^{f})$ requires gluing a stationary null hypersurface $N_{-k}$ for the generalized expansion, subject to specific initial data on $N_{-k}$. This will show that there exists a generalized extremal surface $X$ with zero generalized expansions of two null directions [^8] In the Einstein frame, the surface $X$ where we evaluate the von Neumann entropy is the extremal surface. The extremal condition of $X$ in the Einstein frame will lead $X$ to a generalized extremal surface in the $f(R)$ frame., the Wald entropy of this surface is equal to the Wald entropy of $\mu$. The spacetime is then completed via a CPT reflection across $X$, analogous to the procedure in section 3 (see Figure 3).
@@
Figure 3: This is the full maximizing spacetime in the $f(R)$ frame. The characteristic Cauchy slice that we construct to obtain this geometry consists of $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\cup N^{\prime}_{-l}[\tilde{\mu}$].
@@ 5.3 Construction of the Generalized Extremal Surface and the Outer Entropy in the $f(R)$ Frame
Similarly to section 3.2, we complete our construction by a CPT-reflection across $X$ that takes $v\to-v$, $u\to-u$, $x^{i}\to x^{i}$. Under the CPT transformation, $\Phi$, $f^{\prime}(R)$ and $\chi_{i}$ are even, whereas $\Theta_{u}$, $\Theta_{v}$ and $\partial_{v}f^{\prime}(R)$ are odd. All quantities that are odd under CPT vanish on $X$ by construction. Therefore, the CPT-conjugate data satisfies the requisite junction conditions. The results is a second boundary $\tilde{B}$, $B$ and $\tilde{B}$ are connected by a Cauchy slice with three null segments $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\cup N^{\prime}_{-l}[\tilde{\mu}$], we are using prime to represent CPT-conjugated submanifolds, see figure 3. We have now specified all data necessary to uniquely evolve characteristic initial data via the EOM of $f(R)$ gravity. The resulting spacetime $(\mathcal{M}^{\prime},g)$ has a generalized minimal surface $\mu$ with a fixed outer wedge $O_{W}^{f}[\mu]$. This spacetime contains a generalized extremal surface $X$ on the boundary of the inner wedge $I_{W}^{f}[\mu]$, which is homologous to $\mu$ and therefore to the boundary $B$.
Finally, we need to show that this generalized extremal surface $X$ has the least Wald entropy. Given the duality established in (76), $X$ must have a minimal area in the Einstein frame if $X$ is the extremal surface. Therefore, in the $f(R)$ frame, $X$ should have the least Wald entropy if there exist other generalized extremal surfaces. If the constructed spacetime $(\mathcal{M}^{\prime},g^{\prime})$ contains no other generalized extremal surfaces, the proof is complete. If there exists other generalized extremal surfaces we denote it by $X^{\prime}$, define $\overline{X^{\prime}}[\Sigma]$ as the representative of $X^{\prime}$ on Cauchy slice $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\cup N^{\prime}_{-l}[\tilde{\mu}$]. There are in total three cases; If $\overline{X^{\prime}}[\Sigma]$ lies entirely on $N_{-k}$, then, because $N_{-k}$ is stationary with respect to the generalized expansion, we have
@@ 7 Conclusion
A key ingredient of our analysis is a careful treatment of Einstein frames and their holographic interpretation. We first derive holographic description in the Einstein frame, where the theory can be written as Einstein gravity coupled to an auxiliary scalar degree of freedom, and where standard holographic techniques apply directly. Within this setup, we then establish an explicit correspondence between the boundary von Neumann entropy computed in the Einstein frame description and the corresponding entropy in the $f(R)$ frame. Concretely, we track how the reduced density matrix and its entropy transform under the field redefinitions and Weyl rescalings relating the two frames, and we identify the precise map that preserves the physical content of entanglement. Using this entropy dictionary, we translate the outer entropy construction between frames, obtaining a robust outer entropy correspondence that allows us to import the Einstein frame maximization logic and re-express it in purely geometric language in the $f(R)$ frame. This provides the conceptual bridge needed to equate the coarse-grained outer entropy with the Wald functional of the generalized marginally trapped surface.
@@
Building on this foundation, we proceed to derive the outer entropy directly in $f(R)$ gravity. We prove a focusing theorem for an appropriately defined generalized expansion, which serves as a higher-curvature generalization of the Raychaudhuri equation under the usual null energy conditions. With this theorem in hand, we construct a stationary null hypersurface characterized by vanishing generalized expansion, and we formulate the necessary junction conditions for consistently stitching geometries across a hypersurface within $f(R)$ gravity. These tools allow us to explicitly construct a bulk spacetime whose boundary von Neumann entropy saturates the upper bound implied by the outer entropy maximization—i.e., we realize the extremal configuration achieving the coarse-grained maximum compatible with the fixed exterior. Finally, using the $f(R)$ focusing theorem, we show that the simple entropy defined on the boundary is indeed the holographic dual of the outer entropy in $f(R)$ gravity, thereby extending the outer entropy and simple entropy correspondence to this higher-curvature setting and show the second law of the two entropy which agree with the dynamical black hole entropy Hollands:2024vbe .
@@
Looking forward, our results suggest several promising future directions. Since our construction of the coarse graining entropy is non-perturbative in classical gravity (at least to all orders in perturbation theory around equilibrium), it suggests that the non-perturbative construction of the dynamical black hole entropy is highly related to the generalized marginally trapped surface (generalized minimar surface). In a recent paper Ashtekar:2025qqa it was suggested that quasi-local horizon (QLH) is the right geometric concept for far from equilibrium black hole thermodynamics in Einstein gravity. We hope that similar concepts will appear in general higher curvature gravity. Recently there are some discussions about higher curvature focusing theorem near first order dynamic perturbative killing horizon Yan:2024gbz ; Yan:2025vdp , our construction of a non-perturbative focusing theorem in $f(R)$ gravity suggests similar theorem may exists in more general higher curvature gravity, for example, in $f(\text{Riemann})$ gravity. This will help us define HRT surface in general higher curvature gravity, and explain why the Wall entropy and Dong entropy Dong:2013qoa ; Wall:2015raa are the same but with different physical original.
@@ Acknowledgements.
We would like to thank Antony J. Speranza, Zihan Yan and Jiyu Cheng for useful discussions. In particular, we are greatly grateful to Antony J. Speranza for very patient discussions and useful comments on the draft. We are also grateful for the beautiful environment and academic atmosphere of University of Amsterdam, where this work was completed.
@@ Appendix A Cases of $f^{\prime\prime}(R)=0$
Since $g^{uu}=g^{ui}=g^{vi}=0$ on $N_{-k}$, we only need to consider $T_{uv}$ and $T_{ij}$. Imposing the initial free data conditions (129) and (128) yields $\partial_{v}T=0$. Therefore, this will imply $\partial_{v}R|_{V_{I^{j}_{R}}}=0$. In conclusion, even if there exist subregions of $N_{-k}$ where $f^{\prime\prime}(R)=0$, the result $\partial_{v}R=0$ remains correct.